OpenStudy (anonymous):

The radioactive substance strontium-90 has a half-life of 28 years. In other words, it takes 28 years for half of a given quantity of strontium-90 to decay to a non-radioactive substance. The amount of radioactive strontium-90 still present after t years is modeled by the expression 80*2^-t/28 grams. Evaluate the expression for t = 0, t = 28 and t = 56. What does each value of the expression represent?

OpenStudy (anonymous):

calculator exercise

OpenStudy (anonymous):

well not really \(t=0\) gives you 80, the initial amount

OpenStudy (anonymous):

then \(\frac{28}{28}=1\) and in one half life you have half of what you started with so you get \(80\times (\frac{1}{2})^{\frac{28}{28}}=80\times \frac{1}{2}=40\)